# Max-Rank: Efficient Multiple Testing for Conformal Prediction

### v3

@@ Appendix B ALGORITHMIC DETAILS
2:Output: Adjusted p-values $(P^{\text{adj}}_{1},\dots,P^{\text{adj}}_{m})\in\mathbb{R}^{m}$ with FWER control at level $\alpha$.
@@
5:Compute adjusted p-value by comparing observed statistics to the maximum statistics under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{t}_{\max,i}\geq T_{k}]$.
@@
6:Compute adjusted p-value by comparing observed p-values to the minimum p-values under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{p}_{\min,i}\leq P_{k}]$.
@@
The algorithmic descriptions for Westfall and Young (1993)’s max-T and min-P multiple testing corrections are given in Algo. 3 and Algo. 4 respectively. The main paper contains the algorithms for split conformal prediction in the general case (Algo. 1) and our correction max-rank (Algo. 2).
@@ C.1 FDR control
where the expectation is taken over the $m$ hypotheses. FDR control is appropriate in settings in which a limited proportion of false positives can be tolerated in exchange for a greater chance at significant results, such as in genomics studies (Han et al., 2009). Commonly known methods for FDR control include the Benjamini-Hochberg (Benjamini and Hochberg, 1995), Benjamini-Yekutieli (Benjamini and Yekutieli, 2001) and Simes corrections (Simes, 1986) [^6] Simes was originally designed for FWER control, but then adapted to FDR control.. In contrast to classical corrections with FWER control which assume independence, FDR-controlling procedures are generally more robust to dependencies between tests.
@@ D.1 Simulation study: FWER control and sample size
Figure 4: We examine global adjusted significance levels $\hat{\alpha}$ against the desired FWER control level $\alpha=0.05$ for varying correlations $\rho$, number of tests $m$ and sample size $n$. Empirical FWER control rates are directly related to the magnitudes of $m$ and $n$, suggesting that larger number of parallel tests require larger sample sizes. We also see that Bonferroni is consistently more conservative than max-rank across all combinations.
[…]
The observed interactions between quantities seem consistent and suggest a meaningful relationship that warrants future theoretical investigation, as mentioned in § 7. It would be desirable to quantify the expected FWER control achievable for a given significance level $\alpha$, sample size $n$ and number of tests $m$ in analytical form.
@@ D.2 Multiple testing correction baselines
$C_{\text{Emp}}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\prod_{k=1}^{m}\mathbbm{1}[u_{i,k}\leq u_{k}],$ (21)
where $\mathbbm{1}[\cdot]$ is the indicator function, $u_{k}$ is a component of $\mathbf{u}$ as above (equivalent to $1-\alpha$), and $u_{i,k}$ is the cumulative probability of the $i$-th sample under its marginal c.d.f. in the $k$-th dimension (estimated empirically), i.e., $u_{i,k}=\hat{F}_{X_{k}}(x_{i,k})$. We can think of the copula as empirically estimating the fraction of samples jointly satisfying the condition $\mathbbm{1}[u_{i,k}\leq u_{k}]$ in all dimensions, since only in such cases does the product evaluate to one. As there is no analytical solution (Messoudi et al., 2021) we have to search through a grid of possible adjusted significance levels $\hat{\alpha}\in(0,1)$, evaluate them under the copula, and select the largest one satisfying target coverage $C_{\text{Emp}}(1-\hat{\alpha},\dots,1-\hat{\alpha})\geq(1-\alpha)$, thus directly addressing the problem in Eq. 6 (note that $\min q_{*}\Leftrightarrow\max\hat{\alpha}$).
@@ D.3 Conformal prediction metrics: empirical coverage and prediction set size
$\text{Coverage}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\mathbbm{1}\left[\bigcap_{k=1}^{m}(Y_{j,k}\in\hat{C}(X_{j,k}))\right],$ (22)
where $\mathbbm{1}[\cdot]$ is the indicator function. We note that the empirical coverage is a random quantity parametrized by a coverage distribution, and can slightly deviate from target coverage $(1-\alpha)$ based on factors such as the size of the calibration data $|\mathcal{D}_{cal}|$ (Vovk et al., 2005). For efficiency, we compute the average prediction set size across samples and test dimensions as
@@ D.5 Conformal object detection
Table 2: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on BDD100k for individual classes and the class average. Results are averaged across 100 random trials. max-rank consistently provides strong results also on a per-class basis (best results, second best). We have varying object counts per class, equivalent to the number of per-class calibration samples $n$. By class from top to bottom: 44’960 (avg.), 35’748, 2032, 280’277, 883, 5151, 10’975, and 21’950.

### v4

@@ Appendix B ALGORITHMIC DETAILS
The algorithmic descriptions for Westfall and Young (1993)’s max-T and min-P multiple testing corrections are given in Algo. 3 and Algo. 4 respectively. The main paper contains the algorithms for split conformal prediction in the general case (Algo. 1) and our correction max-rank (Algo. 2). A simple implementation of max-rank takes only a few lines of code, and we provide an example in Listing 1. In this example we simulate perfect correlation ($\rho=1.0$) and observe how max-rank exploits the rank ordering and therefore does not adjust the initial significance level (i.e., $\hat{\alpha}=\alpha=0.1$) which returns the (identical) optimal quantile in each test dimension ($m=5$). In contrast, Bonferroni would adjust to $\hat{\alpha}_{\text{Bonf}}=\alpha/m=0.02$ in each dimension. For $\rho\approx 0$ the same quantiles as Bonferroni are recovered, whereas $\rho\in(0,1)$ provides tighter quantiles (with varying degree).
@@
5:Compute adjusted p-value by comparing observed statistics to the maximum statistics under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}[T_{k}\leq\mathbf{t}_{\max,i}]$.
@@
6:Compute adjusted p-value by comparing observed p-values to the minimum p-values under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}[P_{k}\geq\mathbf{p}_{\min,i}]$.
@@
Listing 1: Exemplary Python code for the max-rank correction.
import numpy as np
alpha = 0.1; n = 1000; m = 5; corr = 1.0 # params
# simulate correlated score matrix
cov = np.full((m, m), corr)
np.fill_diagonal(cov, 1)
S = np.random.multivariate_normal(np.zeros(m), cov, n)
# max-rank correction
R = np.argsort(np.argsort(S, axis=0), axis=0) # column-wise rank matrix
r_max = np.max(R, axis=1) # row-wise max-rank vec(r)_max
q = np.ceil((1 - alpha) * (n + 1)) / n # conformal target coverage level
rank_q = np.quantile(r_max, q, axis=0, method="higher") # quantile r_max
adj_quantile = np.sort(S, axis=0)[rank_q] # final corrected quantiles
print(f"Adjusted significance level: {(1 - rank_q / n):.2f}") # 0.10
print(f"Adjusted quantiles: {adj_quantile}") # e.g. [1.32, 1.32, 1.32, 1.32, 1.32]
@@ Appendix C ADDITIONAL BACKGROUND
We provide additional details on FDR control (§ C.1) and the Bonferroni correction (§ C.2).
@@ C.1 FDR control
where the expectation is taken over the $m$ hypotheses. FDR control is appropriate in settings in which a limited proportion of false positives can be tolerated in exchange for a greater chance at significant results, such as in genomics studies (Han et al., 2009). Commonly known methods for FDR control include the Benjamini-Hochberg (Benjamini and Hochberg, 1995), Benjamini-Yekutieli (Benjamini and Yekutieli, 2001) and Simes corrections (Simes, 1986) [^5] Simes was originally designed for FWER control, but then adapted to FDR control.. In contrast to classical corrections with FWER control which assume independence, FDR-controlling procedures are generally more robust to dependencies between tests. Beyond well-known FWER and FDR control, extensions such as $k$-FWER (Romano and Wolf, 2007; Janson and Su, 2016) – which bounds the probability of committing more than $k$ Type-I errors (recovering FWER for $k=1$) – and false discovery exceedance (Genovese and Wasserman, 2004) – which controls tail probabilities of the false discovery proportion (recovering FDR for the expectation) – have also been explored.
@@ Appendix D ADDITIONAL EXPERIMENT DETAILS
We give additional experiment details for our simulation study (§ D.1), multiple testing baselines (§ D.2), metrics (§ D.3), and conformal experiments (§ D.4, § D.5).
@@ D.1 Simulation study: FWER control and sample size
Figure 4: We examine global adjusted significance levels $\hat{\alpha}$ against the desired FWER control level $\alpha=0.05$ for varying correlations $\rho$, number of tests $m$ and sample size $n$ (shading denotes std. deviation across 10 random trials). Empirical FWER control rates are directly related to the magnitudes of $m$ and $n$, suggesting that larger number of parallel tests require larger sample sizes. We also see that Bonferroni is consistently more conservative than max-rank across all combinations.
Since max-rank connects to Westfall and Young (1993)’s corrections as a resampling-based method, we anticipate that its data-driven nature requires an adequate ratio of tests counts $m$ to sample sizes $n$ per test (equivalent to the number of resamples or permutations) in order to empirically provide FWER control. Thus, we verify in Fig. 4 adjusted significance levels for max-rank and Bonferroni across varying combinations of correlation $\rho$, test counts $m$, and samples $n$. Specifically, we consider $\rho\in\{0,0.25,0.5,0.75,1.0\}$, $m\in[1,1000]$, and $n\in\{100,500,1000\}$. In particular, this also includes high-dimensional settings where $m>n$.
Our results confirm the anticipated dependencies. Specifically, empirical FWER control is not guaranteed when the ratio of samples to test dimensions is too low, including the settings where $m>n$. These violations affect both max-rank and Bonferroni, albeit our correction tends to violate sooner since it operates closer to target level. However, FWER control rates improve substantially as both sample size and correlation increases. Intuitively, higher correlation effectively reduces the amount of independent rank order information across tests that needs to be accounted for. Overall, we observe that empirical FWER control is approximately maintained when $n\geq 100\cdot m$ in the least favorable setting ($\rho=0$). However, there are cases where this rate can be exceeded (e.g., $n=100$ and $m=5$ holds) or fall short (e.g., $n=1000$ and $m=10$ undercorrects slightly). Finally, throughout all considered combinations Bonferroni is always more conservative than max-rank (in line with Prop. 2), and its disadvantage to max-rank increases substantially as the correlation goes up.
The observed sample requirements seem reasonable given that a reliable empirical null distribution must be established for each test dimension, and are somewhat in the range of calibration set sizes commonly used in conformal prediction. For instance, Angelopoulos et al. (2023) recommend at least 1000 calibration samples for a robust conformal procedure even in a single dimension, while Messoudi et al. (2021) (see Table 2 in their paper) consider varying sample requirements for multi-target regression such as $(n,m)\in\{(2100,3),(24160,4),(910,8)\}$. The impact of overly small sample sizes on results is evident both in their work (e.g., Fig. 3 in their paper) and ours (Table 1, for the scpf dataset).
In general, the observed interactions between quantities seem consistent and suggest a meaningful relationship that warrants future theoretical investigation, as mentioned in § 7. It would be desirable to quantify the expected FWER control achievable for a given significance level $\alpha$, sample size $n$ and number of tests $m$ in analytical form.
@@ D.2 Multiple testing correction baselines
$C_{\text{Emp}}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\prod_{k=1}^{m}\mathds{1}[u_{i,k}\leq u_{k}],$ (21)
where $\mathds{1}[\cdot]$ is the indicator function, $u_{k}$ is a component of $\mathbf{u}$ as above (equivalent to $1-\alpha$), and $u_{i,k}$ is the cumulative probability of the $i$-th sample under its marginal c.d.f. in the $k$-th dimension (estimated empirically), i.e., $u_{i,k}=\hat{F}_{X_{k}}(x_{i,k})$. We can think of the copula as empirically estimating the fraction of samples jointly satisfying the condition $\mathds{1}[u_{i,k}\leq u_{k}]$ in all dimensions, since only in such cases does the product evaluate to one. As there is no analytical solution (Messoudi et al., 2021) we have to search through a grid of possible adjusted (global) significance levels $\hat{\alpha}\in(0,1)$, evaluate them under the copula, and select the largest one satisfying target coverage $C_{\text{Emp}}(1-\hat{\alpha},\dots,1-\hat{\alpha})\geq(1-\alpha)$, thus directly addressing the problem in Eq. 6 (note that $\min q_{*}\Leftrightarrow\max\hat{\alpha}$).
@@ D.3 Conformal prediction metrics: empirical coverage and prediction set size
$\text{Coverage}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\mathds{1}\left[\bigcap_{k=1}^{m}(Y_{j,k}\in\hat{C}(X_{j,k}))\right],$ (22)
where $\mathds{1}[\cdot]$ is the indicator function. We note that the empirical coverage is a random quantity parametrized by a coverage distribution, and can slightly deviate from target coverage $(1-\alpha)$ based on factors such as the size of the calibration data $|\mathcal{D}_{cal}|$ (Vovk et al., 2005). For efficiency, we compute the average prediction set size across samples and test dimensions as
@@ D.5 Conformal object detection
Table 3: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on BDD100k for individual classes and the class average. Results are averaged across 100 random trials. max-rank consistently provides strong results also on a per-class basis (best, second best results). We have varying object counts per class, equivalent to the number of per-class calibration samples $n$. From top to bottom by class: 44’960 (Class avg.), 35’748, 2032, 280’277, 883, 5151, 10’975, and 21’950.